# DeepInteraction++: Multi-Modality Interaction for Autonomous Driving

### v2

@@ V-E Extension to the end-to-end planning
We train the e2e framework with the same settings as the detection task, except for a batch size of 1. As for metrics, $minADE$, $minFDE$, and $MR$ across six prediction modes are employed for the evaluation of prediction performance. In the context of the planning task, we adopt ego-trajectory displacement error (L2) and Collision Rate as the measure.
@@
Benefiting from the multi-modal representation and interactive decoding, our e2e extension of DeepInteraction++ achieves better perception and prediction performance compared to VAD [47], as shown in Table XI. Moreover, we report the planning results in Table X, which demonstrates that Deep Interaction++ remarkably surpasses existing planning-oriented methods on most evaluation metrics. Besides providing a more accurate planning trajectory, DeepInteraction++ can achieve a lower collision rate by resorting to more precise and comprehensive perception and prediction for traffic participants. Furthermore, we also implement an end-to-end framework based on the original version DeepInteraction, which takes the sparse points as a medium for representation interaction. In comparison, the DeepInteraction++ can preserve more road elements from images thorough deformable attention and dense polar interaction, achieving superior performance across all metrics.
@@ VI Conclusion
In this work, we have presented a novel multi-modality interaction approach for exploring both the intrinsic multi-modal complementary nature and their respective characteristics in autonomous driving. This key idea is to maintain two modality-specific representations and establish interactions between them for both representation learning and predictive decoding. This strategy is designed particularly to resolve the fundamental limitation of existing unilateral fusion approaches that image representation is insufficiently exploited due to their auxiliary-source role treatment. Extensive experiments demonstrate our approach yields state-of-the-art performances on the highly-competitive nuScenes benchmark, across both 3D object detection and more other end-to-end autonomous driving tasks.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62106050 and 62376060), the Natural Science Foundation of Shanghai (Grant No. 22ZR1407500), and Lingang Laboratory (Grant No. LG-QS-202202-07).

### v3

@@ V-E Extension to the end-to-end planning
We train the e2e framework with the same settings as the detection task, except for a batch size of 1. As for metrics, $minADE$, $minFDE$, and $MR$ across six prediction modes are employed for the evaluation of prediction performance, while ego-trajectory displacement error (L2) and Collision Rate are adopted in the planning task.
@@
Benefiting from the multi-modal representation and interactive decoding, our e2e extension of DeepInteraction++ achieves better perception and prediction performance compared to VAD [47], as shown in Table XIV. Moreover, we report the planning results in Table XIII, which demonstrates that DeepInteraction++ remarkably surpasses existing planning-oriented methods on most evaluation metrics. Besides providing a more accurate planning trajectory, DeepInteraction++ can achieve a lower collision rate by resorting to more precise and comprehensive perception and prediction for traffic participants. Furthermore, we also implement an end-to-end framework based on the original DeepInteraction, which takes the sparse points as a medium for representation interaction. In comparison, the DeepInteraction++ can preserve more road elements from images thorough deformable attention and dense polar interaction, achieving superior performance across all metrics.
@@
V-F Failure cases and discussions
In Figure 9, we present several failure cases to provide a more comprehensive perspective on the limitations of the proposed framework and shed light on potential challenges may face in practice. For scene perception, our explicit LiDAR-guided 3D mapping makes the model susceptible to misalignment or sensor failures. For instance, if an object lacks LiDAR signals, it may be missed by the detector. While the integration of learning-based polar ray attention helps mitigate this issue to some extent, it still occurs in certain cases, as illustrated in the top plot of Figure 9. For planning tasks, although multi-sensor fusion provides richer scene information, challenges such as map segmentation and motion prediction remain not fully resolved within existing frameworks. As a result, this can lead to unreasonable planning trajectories, as shown in the bottom plot of Figure 9.
@@ VI Conclusion
In this work, we have presented a novel multi-modality interaction approach for exploring both the intrinsic multi-modal complementary nature and their respective characteristics in autonomous driving. This key idea is to maintain two modality-specific representations and establish interactions between them for both representation learning and predictive decoding. This strategy is designed particularly to resolve the fundamental limitation of existing unilateral fusion approaches that image representation is insufficiently exploited due to their auxiliary-source role treatment. Extensive experiments demonstrate our approach yields state-of-the-art performances on the highly-competitive nuScenes benchmark, across both 3D object detection and end-to-end autonomous driving tasks.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62376060).